# Quadratic Speed-up in Infinite Variance Quantum Monte Carlo

### 15 Jan 2024 v1

@@ Abstract
In this study, we give an extension of Montanaro’s [Mon15] quantum Monte Carlo method, tailored for computing expected values of random variables that exhibit infinite variance. This addresses a challenge in analyzing heavy-tailed distributions, which are commonly encountered in various scientific and engineering fields. Our quantum algorithm efficiently estimates means for variables with a finite $(1+\delta)^{\text{th}}$ moment, where $\delta$ lies between 0 and 1. It provides a nearly quadratic speedup over the classical Monte Carlo method in both the accuracy parameter $\epsilon$ and the specified moment of the distribution. We establish both classical and quantum lower bounds, showcasing the near-optimal efficiency of our algorithm among quantum methods. Our work focuses not on creating new algorithms, but on analyzing the execution of existing algorithms with available additional information about the random variable. Additionally, we categorize these scenarios and demonstrate a hierarchy in the types of supplementary information that can be provided.
@@ 1 Introduction
Heavy-tailed phenomena can explain empirical findings in internet traffic, earth sciences, and other areas, leading to long-range dependence and persistent correlations in time series data. In the context of internet traffic, a server experiences packet requests of users navigating the internet at a high frequency, probably in the order of seconds or even faster, whereas users consume information requested in the order of many minutes or perhaps even hours. This creates on-and-off interactions between the user and the server which relative to the server’s time scale of operation are better modeled as very heavy-tailed random times [TTW95]. These models were pioneered by Mandelbrot [Man63] and have been studied in various fields; see [PT17] and [Wat19].
Heavy-tailed distributions (potentially with infinite variance), however, can also pose challenges for statistical estimation, leading to inaccurate predictions and risk assessments. Consequently, accurately computing expected values for such distributions is a critical practical necessity.
Quantum computing offers potential speed-ups for various computational problems, including Monte Carlo methods. The work by [Mon15] introduced a quantum Monte Carlo algorithm with quadratic speed-up over classical methods for estimating the expected value of a random variable with finite variance. The algorithm relies on phase estimation [BHT98, BHMT02] which is in turn grounded in Grover’s algorithm [Gro96]. It has also been applied in quantum Markov Chain Monte Carlo [${}^{+}$22] and computational finance [${}^{+}$20].
In this study, we extend Montanaro’s method to estimate expected values for random variables with infinite variance but finite $(1+\delta)^{th}$ moment, where $\delta\in(0,1]$. Our algorithm exhibits a quadratic speed-up in relation to both the accuracy parameter $\epsilon$ and the $(1+\delta)^{th}$ moment of the distribution. It achieves $\epsilon$-Mean absolute error (MAE), a stronger metric than the ‘large probability guarantee’ used by most existing quantum algorithms.
We provide both classical and quantum lower bounds, demonstrating that our quantum algorithm nearly quadratically surpasses all classical algorithms, not just traditional Monte Carlo methods, and is nearly optimal (within polylogarithmic factors) among all quantum algorithms with respect to both $\epsilon$ and the $(1+\delta)^{th}$ moment. Our lower bounds are established by reducing Grover-type search problems to our mean estimation problems.
Montanaro’s algorithm relies on phase estimation, as described by [BHT98, BHMT02], that in turn are deeply grounded in Grover’s algorithm [Gro96]. This algorithm has also been applied in various other algorithms and fields, including quantum Markov Chain Monte Carlo [${}^{+}$22] and computational finance [${}^{+}$20]. In 2023, Kothari and O’Donnell [KO23] further improved Montararo’s work, which we utilize as a black box in our study.
@@ 2 Quantum Monte Carlo algorithm
The canonical problem in Monte Carlo is that of computing $\mathbb{E}[X]$. Its current quantized versions [Mon15, Ham21, KO23] focus on estimating the expectation of a random variable with finite variance. We show how to use similar ideas to these results (and in particular to [Mon15, KO23]), when the random variable has an infinite variance. We do need some hypothesis, however, and we will assume, that there is a $0<\delta\leq 1$ and a known $C_{1+\delta}$ such that $\mathbb{E}[\lvert v({\mathcal{A}})\rvert^{1+\delta}]\leq C_{1+\delta}<\infty$.
@@
There is a procedure that, given any (classical or quantum) algorithm $\mathcal{A}$ which outputs a random variable $v(\mathcal{A})$ satisfying $\mathbb{E}[\lvert v({\mathcal{A}})\rvert^{1+\delta}]\leq C_{1+\delta}<\infty$ for some $\delta\in(0,1]$, and an upper bound $C$ on $C_{1+\delta}$, constructs a quantum algorithm which estimates $\mathbb{E}[v({\mathcal{A}})]$ up to additive error $\epsilon$ with success probability at least $4/5$ by using ${\mathcal{O}}\left(C^{1/2\delta}\epsilon^{-(1+\delta)/2\delta}\right)$ samples (i.e. applications of a quantum-adapted version of ${\mathcal{A}}$, as in [KO23]). The hidden factor within ${\mathcal{O}}$ depends on $\delta$.
@@
The result above establishes a quadratic speedup when compared to the classical or ’vanilla’ Monte Carlo algorithm, which is optimal within a constant factor, and operates by simply outputting the average of the samples it generated.
@@ 3 The Model
To answer the first question: We adopt the model presented in [KO23]. Classical sampling theory is based on an oracle model, in which a ”button” is pushed to obtain a random sample of the variable. We can think of any random variable of interest as $X(u)$, where $u\in[0,1]$. When the button is pushed, the oracle (or ’automaton’) generates a uniformly random $u\in[0,1]$ and provides us with the corresponding value $X(u)$. The complexity of an algorithm is measured by the number of times the button is pushed.
@@
The Code of Kothari and O’Donnell [KO23]. We call ${\cal A}$ ’the code,’ which is either an input-less randomized circuit or a quantum circuit with input $|0^{n}\rangle$ and a final measurement of $k$ qubits that returns the random variable $X=v({\cal A})\in\mathbb{R}$. While Montanaro also assumes the availability of the code for ${{\mathcal{A}}}$ and ${{\mathcal{A}}}^{-1}$, he modifies it less than [KO23]. Kothari and O’Donnell post-process ${\cal A}$ to turn it into the phase shift operator $e^{-2i\arctan v({\cal A})}$. With this definition Kothari and O’Donnell prove:
@@ 3.1 Infinite Random Variables
When we have an infinite $X$, there are different ways to think about it in an algorithmic framework.
An infinite ${\cal A}$. There exist classical algorithms representing infinite random variables. Consider an experiment in which we perform consecutive coin flips. After the first instance of ’heads’ appears following $k$ occurrences of ’tails,’ we output $2^{\alpha k}$ (where $\alpha<1$). The output, $v({\mathcal{A}})$, possesses an infinite variance as long as $\alpha>1/2$. However, it has a finite mean and a finite $(1+\delta)^{\rm th}$ moment for any $\delta<\frac{1}{\alpha}-1$. In attempts to quantize such a model, we encounter challenges due to the variable-time nature of ${\cal A}$.
@@
A finite ${\cal A}$ with a large $\mathbb{E}[\lvert v({\mathcal{A}})\rvert^{2}]$ but a small $\mathbb{E}[\lvert v({\mathcal{A}})\rvert^{1+\delta}]$. Here our attitude is that the expression ’heavy tail’ should be taken with a grain of salt, since in reality no random variable is infinite, and what in practice ’heavy tail’ really means is that $\mathbb{E}[X^{2}]$ is very large, but $\mathbb{E}[X^{1+\delta}]$ is relatively small. It is sufficient to say that we estimate the running time of the quantization in terms of the parameter $\mathbb{E}[X^{1+\delta}]$, regardless of $\mathbb{E}[X^{2}]$.
The type of sampling problem we address in this article involves a given target additive error $\epsilon$ for the mean estimation problem. We must assume that the algorithm initially receives an upper bound $C$ on $\mathbb{E}[X^{1+\delta}]$ as in Theorem 1. Theorem 1 could not hold with any finite bound on the sample size if $C$ were not given as part of the input. Our algorithm then proceeds by determining an appropriate cutoff value, depending on both $\epsilon$ and $\mathbb{E}[X^{1+\delta}]$. After the application of this cutoff, the truncated variable becomes finite, rendering the distinction between the three aforementioned methods of thinking about an infinite variable irrelevant.
@@ 4 Background on moments
Therefore, one can choose $n={\mathcal{O}}(\epsilon^{-p/(p-1)})={\mathcal{O}}(\epsilon^{-(1+\delta)/\delta})$ to ensure $\epsilon$-error.
@@ 6 Lower bounds
The classical portion of Theorem 6 seems folklore, at least we could not identify a source, but no researchers in the area would doubt it is true.

### 07 Mar 2024 v2

@@ Abstract
In this study, we give an extension of Montanaro’s [Mon15] quantum Monte Carlo method, tailored for computing expected values of random variables that exhibit infinite variance. This addresses a challenge in analyzing heavy-tailed distributions, which are commonly encountered in various scientific and engineering fields. Our quantum algorithm efficiently estimates means for variables with a finite $(1+\delta)^{\text{th}}$ moment, where $\delta$ lies between 0 and 1. It provides a quadratic speedup over the classical Monte Carlo method in both the accuracy parameter $\epsilon$ and the specified moment of the distribution. We establish both classical and quantum lower bounds, showcasing the near-optimal efficiency of our algorithm among quantum methods. Our work focuses not on creating new algorithms, but on analyzing the execution of existing algorithms with available additional information about the random variable. Additionally, we categorize these scenarios and demonstrate a hierarchy in the types of supplementary information that can be provided.
@@ 1 Introduction
Heavy-tailed phenomena can explain empirical findings in internet traffic, earth sciences, and other areas, leading to long-range dependence and persistent correlations in time series data. In the context of internet traffic, a server experiences packet requests of users navigating the internet at a high frequency, probably in the order of seconds or even faster, whereas users consume information requested in the order of many minutes or perhaps even hours. This creates on-and-off interactions between the user and the server which relative to the server’s time scale of operation are better modeled as very heavy-tailed random times [TTW95]. These models were pioneered by Mandelbrot [Man63] and have been studied in various fields; see [PT17] and [Wat19]. Heavy-tailed distributions (potentially with infinite variance), however, can also pose challenges for statistical estimation, leading to inaccurate predictions and risk assessments. Consequently, accurately computing expected values for such distributions is a critical practical necessity.
Quantum computing offers potential speed-ups for various computational problems, including Monte Carlo methods. The work by [Mon15] introduced a quantum Monte Carlo algorithm with quadratic speed-up over classical methods for estimating the expected value of a random variable with finite variance, which generalizes the algorithm of [Hei03] for mean estimation under the uniform distribution. Montanaro’s algorithm relies on phase estimation [BHT98, BHMT02] which is in turn grounded in Grover’s algorithm [Gro96]. It has also been applied in quantum Markov Chain Monte Carlo [${}^{+}$22] and computational finance [${}^{+}$20]. Hamoudi and Magniez [HM19] improve Montanaro’s algorithm in terms of relative error, assuming that an upper and lower bound for the expected value is known to the users. The thesis of Hamoudi [Ham21] also improves Montanaro’s algorithm regarding additive errors, eliminating the need for a pre-determined upper bound on the variance. Recently, Kothari and O’Donnell [KO23] further improved existing works, which we utilize as a black box in our study.
In this study, we extend existing methods to estimate expected values for random variables with infinite variance but finite $(1+\delta)^{th}$ moment, where $\delta\in(0,1]$. Our algorithm exhibits a quadratic speed-up in relation to both the accuracy parameter $\epsilon$ and the $(1+\delta)^{th}$ moment of the distribution. It achieves $\epsilon$-Mean absolute error (MAE), a stronger metric than the ‘large probability guarantee’ used by most existing quantum algorithms.
We provide both classical and quantum lower bounds, demonstrating that our quantum algorithm nearly quadratically surpasses all classical algorithms, not just traditional Monte Carlo methods, and is nearly optimal among all quantum algorithms with respect to both $\epsilon$ and the $(1+\delta)^{th}$ moment. Our lower bounds are established by reducing Grover-type search problems to our mean estimation problems.
@@ 2 Quantum Monte Carlo algorithm
The canonical problem in Monte Carlo is that of computing $\mathbb{E}[X]$. Its current quantized versions [Mon15, Ham21, KO23] focus on estimating the expectation of a random variable with finite variance. We show how to use similar ideas to these results (and in particular to [Mon15, KO23]), when the random variable has an infinite variance. We do need some hypothesis: we will assume that there is a $0<\delta\leq 1$ and a known $C$ such that $\mathbb{E}[\lvert v({\mathcal{A}})\rvert^{1+\delta}]:=C_{1+\delta}\leq C<\infty$.
@@
There is a procedure that, given any (classical or quantum) algorithm $\mathcal{A}$ which outputs a random variable $v(\mathcal{A})$ satisfying $\mathbb{E}[\lvert v({\mathcal{A}})\rvert^{1+\delta}]=C_{1+\delta}<\infty$ for some $\delta\in(0,1]$, and an upper bound $C$ on $C_{1+\delta}$, constructs a quantum algorithm which estimates $\mathbb{E}[v({\mathcal{A}})]$ up to additive error $\epsilon$ with success probability at least $4/5$ by using ${\mathcal{O}}\left(C^{1/2\delta}\epsilon^{-(1+\delta)/2\delta}\right)$ samples (i.e. applications of a quantum-adapted version of ${\mathcal{A}}$, as in [KO23]). The hidden factor within ${\mathcal{O}}$ depends on $\delta$.
@@
The result above establishes a quadratic speedup when compared to the classical or ‘vanilla’ Monte Carlo algorithm, which is optimal within a constant factor, and operates by simply outputting the average of the samples it generated.
@@ 3 The Model
To answer the first question: We adopt the model presented in [KO23]. Classical sampling theory is based on an oracle model, in which a ”button” is pushed to obtain a random sample of the variable. We can think of any random variable of interest as $X(u)$, where $u\in[0,1]$. When the button is pushed, the oracle (or ‘automaton’) generates a uniformly random $u\in[0,1]$ and provides us with the corresponding value $X(u)$. The complexity of an algorithm is measured by the number of times the button is pushed.
@@
The Code of Kothari and O’Donnell [KO23]. We call ${\cal A}$ ‘the code,’ which is either an input-less randomized circuit or a quantum circuit with input $|0^{n}\rangle$ and a final measurement of $k$ qubits that returns the random variable $X=v({\cal A})\in\mathbb{R}$. While Montanaro also assumes the availability of the code for ${{\mathcal{A}}}$ and ${{\mathcal{A}}}^{-1}$, he modifies it less than [KO23]. Kothari and O’Donnell post-process ${\cal A}$ to turn it into the phase shift operator $e^{-2i\arctan v({\cal A})}$. With this definition Kothari and O’Donnell prove:
@@ 3.1 Infinite Random Variables
When we have a random variable $X$ that does not have a bounded support, there are different ways to think about it in an algorithmic framework.
An infinite ${\cal A}$. There exist classical algorithms representing infinite random variables. Consider an experiment in which we perform consecutive coin flips. After the first instance of ‘heads’ appears following $k$ occurrences of ‘tails,’ we output $2^{\alpha k}$ (where $\alpha<1$). The output, $v({\mathcal{A}})$, possesses an infinite variance as long as $\alpha>1/2$. However, it has a finite mean and a finite $(1+\delta)^{\rm th}$ moment for any $\delta<\frac{1}{\alpha}-1$. In attempts to quantize such a model, we encounter challenges due to the variable-time nature of ${\cal A}$.
@@
A finite ${\cal A}$ with a large $\mathbb{E}[\lvert v({\mathcal{A}})\rvert^{2}]$ but a small $\mathbb{E}[\lvert v({\mathcal{A}})\rvert^{1+\delta}]$. Here our attitude is that the expression ‘heavy tail’ should be taken with a grain of salt, since in reality no random variable is infinite, and what in practice ‘heavy tail’ really means is that $\mathbb{E}[X^{2}]$ is very large, but $\mathbb{E}[X^{1+\delta}]$ is relatively small. It is sufficient to say that we estimate the running time of the quantization in terms of the parameter $\mathbb{E}[\lvert X\rvert^{1+\delta}]$, regardless of $\mathbb{E}[X^{2}]$.
The type of sampling problem we address in this article involves a given target additive error $\epsilon$ for the mean estimation problem. We must assume that the algorithm initially receives an upper bound $C$ on $\mathbb{E}[\lvert X\rvert^{1+\delta}]$ as in Theorem 1. Theorem 1 could not hold with any finite bound on the sample size if $C$ were not given as part of the input. Our algorithm then proceeds by determining an appropriate cutoff value, depending on both $\epsilon$ and $\mathbb{E}[\lvert X\rvert^{1+\delta}]$. After the application of this cutoff, the truncated variable becomes finite, rendering the distinction between the three aforementioned methods of thinking about an infinite variable irrelevant.
@@ 4 Background on moments
Therefore, one can choose $n={\mathcal{O}}(\epsilon^{-p/(p-1)})={\mathcal{O}}(\epsilon^{-(1+\delta)/\delta})$ to ensure the ‘vanilla’ Monte Carlo estimator has at most $\epsilon$-error.
@@ 6 Lower bounds
The classical portion of Theorem 6 is also proven in Theorem 3.1 of [DLLO16] in a different way.